Please help and answer in a PSEUDO Code format in assembly language! Thanks

PROJECT (a) write a program that will convert the number 23567 => n, where (b) Write an algorithm that will convert a number n, to ne where b , c < 10.
Answer #1

a) 23567=>nb; b=9

     Begin:

               Input: 2356

               Output: nb

               Convert Decimal:

                        initialise i=0, sum=0

                        while input not equal 0

                                sum=sum+(input%10)*7i

                                input=input/10

                                increment i by 1

                         end while

                  return sum

            Convert Base 9:

                            initialise val=0

                            initialise i=0

                            while sum not equal 0

                                    val=(sum%9)*10i +val

                                    sum=sum/9

                                    increment i by 1

                              end while

              return val

      End

     Answer is 23567=11719

b) nb=nc   b,c<10

     Begin:

               Input: nb

               Output: nc

               Convert Decimal:

                        initialise i=0, sum=0

                        while input not equal 0

                                sum=sum+(input%10)*bi

                                input=input/10

                                increment i by 1

                         end while

                  return sum

            Convert Base c:

                            initialise val=0

                            initialise i=0

                            while sum not equal 0

                                    val=(sum%c)*10i +val

                                    sum=sum/c

                                    increment i by 1

                              end while

              return val

      End